OpenStudy (anonymous):

find the value of x x x+20 2x

OpenStudy (anonymous):

Any answer choices ?

OpenStudy (raden):

retype your equation, please ?

OpenStudy (anonymous):

find the value of x. x, x+20, and 2x

OpenStudy (anonymous):

there are are no answer choices have to write it out

OpenStudy (raden):

add up all to get sum 180 x + 2x + x + 20 = 180 solve for x

OpenStudy (anonymous):

how would you do that

OpenStudy (anonymous):

when i solved it i got 7.5. went bck to check my answer i got 50. wtf need help

OpenStudy (raden):

combine the similar terms to simplify that equation : x + 2x + x + 20 = 180 4x + 20 = 180

OpenStudy (raden):

subtracted by 20 from both sides, 4x + 20 = 180 4x + 20 - 20 = 180 - 20 4x = 160

OpenStudy (raden):

then divide by 4 to both sides, 4x = 160 4x/4 = 160/4 x = 40

OpenStudy (anonymous):

omg thanks much. so if i go bck to check it im gonna get 180????